OpenStudy (tiffany_rhodes):

find:

OpenStudy (tiffany_rhodes):

\[\int\limits_{3}^{5}\int\limits_{2}^{6}\]

OpenStudy (tiffany_rhodes):

xye^(x+y) dydx

ganeshie8 (ganeshie8):

Hint : \[e^{x+y} = e^x\cdot e^y\]

OpenStudy (tiffany_rhodes):

hmm, I'm assuming I'm going to have to use IBP?

ganeshie8 (ganeshie8):

yes it will be a trivial IBP

ganeshie8 (ganeshie8):

\[\int xe^x = xe^x - e^x\]

ganeshie8 (ganeshie8):

+C

OpenStudy (tiffany_rhodes):

Ooh, you can use separation of variables, correct?

OpenStudy (tiffany_rhodes):

and just split up the integrals in terms of only x and only y?

ganeshie8 (ganeshie8):

\[\int\limits_{3}^{5}\int\limits_{2}^{6} xye^{x+y}dydx = \left(\int\limits_3^5 xe^x\right)\cdot \left(\int\limits_2^6ye^ydy\right)\] like this ?

OpenStudy (tiffany_rhodes):

yes

ganeshie8 (ganeshie8):

that works but still u need to use IBP

OpenStudy (tiffany_rhodes):

Okay.

OpenStudy (tiffany_rhodes):

got it. Thanks @ganeshie8